### v1

# Unified Software Engineering agent as AI Software Engineer

Abstract.
The growth of Large Language Model (LLM) technology has raised expectations for automated coding. However, software engineering is more than coding and is concerned with activities including maintenance and evolution of a project. In this context, the concept of LLM agents has gained traction, which utilize LLMs as reasoning engines to invoke external tools autonomously. But is an LLM agent the same as an AI software engineer? In this paper, we seek to understand this question by developing a Unified Software Engineering agent or USEagent. Unlike existing work which builds specialized agents for specific software tasks such as testing, debugging, and repair, our goal is to build a unified agent which can orchestrate and handle multiple capabilities. This gives the agent the promise of handling complex scenarios in software development such as fixing an incomplete patch, adding new features, or taking over code written by others. We envision USEagent as the first draft of a future AI Software Engineer which can be a team member in future software development teams involving both AI and humans. To evaluate the efficacy of USEagent, we build a Unified Software Engineering bench (USEbench) comprising of myriad tasks such as coding, testing, and patching. USEbench is a judicious mixture of tasks from existing benchmarks such as SWE-bench, SWT-bench, and REPOCOD. In an evaluation on USEbench consisting of 1,271 repository-level software engineering tasks, USEagent shows improved efficacy compared to existing general agents such as OpenHands CodeActAgent. For specific tasks such as issue resolution in SWE-bench, USEagent demonstrates efficacy comparable to agents such as AutoCodeRover (which are focused on software maintenance), while maintaining applicability to a broader range of software engineering tasks. There exist gaps in the capabilities of USEagent for certain coding tasks, which provides hints on further developing the AI Software Engineer of the future.
Automated Software Engineering, Agentic Systems, AI for Software Development
*Joint first authors, ordered alphabetically.
[…]
ccs: Software and its engineering Automatic programmingccs: Software and its engineering Software maintenance toolsccs: Software and its engineering Software evolution
1. Introduction
Large language models (LLMs) have shown promise in coding, reasoning, and problem solving. The model capability itself has improved significantly in terms of reasoning tasks, by training the models to always use reasoning paradigms such as chain of thoughts, without needing further prompting, as shown by GPT o1 (OpenAI, 2024b) and its descendant models such as o3. On top of improved foundation models, the software engineering community is distilling known best practices for downstream tasks into amplifiers for LLMs through agentic systems (Yao et al., 2023). An LLM agent for software engineering invokes various interfaces including testing and analysis tools autonomously driven by an LLM reasoning agent.
While agentic systems inspire researchers (Yao et al., 2023; Antoniades et al., 2024), start-ups (Orwall and contributors, 2025; AIDER Team, 2024) and major companies (Wadhwa et al., 2024; Rondon et al., 2025), we currently observe a strong specialization in the emerging technologies among the LLM agents for software. Agentless (Xia et al., 2024) performs program repair, Large Language Monkeys (Brown et al., 2024) produce unit tests, ExecutionAgent (Bouzenia and Pradel, 2024) performs a project setup, and so forth. These specialized agents are reasonable efforts, especially as they require specialized metrics and datasets. However, with many agents comes the need to manage and maintain them as part of the future development environments!
We propose a unified Software Engineering agent (USEagent) representing a consolidated agentic capability for software engineering tasks. Individual SE tasks could be coordinated to result in a more effective ensemble of tasks: better test-generation should benefit reproduction in program repair, review of patches should lead to better generated code, etc. With an integrated software engineering agent, it becomes much more feasible to co-opt it in future development environments (A Roychoudhury and Ray, 2025).
As a foundation, we need a unified dataset of challenging SE tasks to test our unified agentic capabilities, for which we build Unified Software Engineering Bench, or USEbench for short. Recently, the SWE-bench dataset (Jimenez et al., 2024) has been proposed that captures a set of GitHub issues depicted in natural language and requires bug fixes or feature additions in software projects. Thus the SWE-bench dataset is replete with challenges in software maintenance for various real-life software projects. Our proposed benchmark USEbench is a meta-benchmark composing multiple software engineering tasks (such as code generation, program repair, test generation) behind a unified application programming interface (API). Building a unified API is key, since it allows us to start thinking of a unified agentic capability which can handle different software engineering tasks. After constructing USEbench, we expand two popular agentic systems, one from industry and one from academia, AutoCodeRover (Zhang et al., 2024b) and OpenHands CodeActAgent (Wang et al., 2024b, a) to solve the tasks in USEbench. We consider this a natural evolution or progression of the flurry of research proposing manifold different agents for different software engineering tasks. Existing specialized agents usually employ a fixed workflow and approach the given task in a few pre-defined steps. For example, AutoCodeRover tackles program repair tasks using a fixed two-phase workflow consisting of fault localization and patch generation. Now, what does it take to generalize an existing program repair agent, with fixed actions and pre-defined workflow, so that it can act as a unified software engineering agent (USEagent)? For each task or problem in USEbench, the agentic systems are only provided with a description (i.e., a bug report or the documentation of a method to generate) as well as access to a containerized environment of the project. As such, our meta-benchmark reveals inherent challenges regarding task-identification, workflow-configuration and measuring progress (esp. determining end-criteria).
To address these requirements, we equip AutoCodeRover with a Meta-Agent, which is instructed to orchestrate the appropriate agents and construct a workflow on the fly. Over the course of this workflow, the Meta-Agent utilizes available actions to construct and maintain a project-state, constituting a structured consensus memory over the trajectories LLM components. Since OpenHands CodeActAgent is a general agent designed for solving general tasks, we consider it for baseline comparison, with no architectural changes applied.
We report the efficacy of both agentic systems on USEbench including PASS@1 and PASS@5 results. We perform a detailed error analysis as well as a manual inspection of positives, to identify false-positives and cases of over-fitting or memorization. On USEbench consisting of 1271 tasks including program repair, regression testing, code generation, and test generation, USEagent achieves an efficacy of 33.3%, which is higher than the 26.8% efficacy from the state-of-the-art general agent OpenHands CodeActAgent. Specifically, on software maintenance tasks in SWE-bench-verified, USEagent has an efficacy of 45.6% which is similar to the 46.2% efficacy of the specialized AutoCodeRover agent, while being applicable to more types of tasks. On test generation tasks that AutoCodeRover could not be applied to, USEagent achieves 31.8% efficacy, demonstrating its versatility. We also make the USEbench benchmark public, to encourage further research in design of AI software engineers.
2. Unified Software Engineering Benchmark
Benchmark Task Concept Solving Criteria #
Primitive Tasks
SWE-bench-verified Program Repair Adjust a program according to an issue. Test-Suite Pass 500
SWT-bench-Lite Regression Testing Produce a test that asserts a required change, as presented in an issue Patch-Coverage 298
REPOCOD-Lite Code Generation Generate a method body, matching the functionality outlined in documentation Test-Suite Pass 200
REPOTEST-Lite Test Generation Test a given method to reach 100% coverage Code-Coverage 173
Compound Tasks
Partial fix Test Generation
Program Repair
Enrich the SWE-Task with a previously failed, but promising patch Test-Suite Pass 100
Feature Development Code- & Test-Generation Add both tests and code, evaluate against the newly produced code coverage Code-Coverage 46
Table 1. Summary of primitive and compound tasks unified in USEbench
To make progress towards building a unified agent for software engineering (USEagent), we first build a dataset of automated software engineering tasks. There is a rich environment of benchmarks, yet they are somewhat fragmented - they focus on specific types of software engineering tasks. Popular benchmarks encompass natural language issues for software (SWE-bench (Jimenez et al., 2024)), automated program repair (Defects4J (Just et al., 2014)), code generation (REPOCOD (Liang et al., 2024)), test generation (SWT (Mündler et al., 2024)) or documentation generation (CodeNet (Puri et al., 2021)), each accompanied by a metric to identify correct or optimal solutions. Next to their task, benchmarks often vary in scope too: the programming benchmarks often considered in foundational model evaluation sets (such as HumanEval (Chen et al., 2021b) , CodeNet (Puri et al., 2021), BigCode (Anonymous, 2021), EvalPlus (Liu et al., 2024)) cover method-snippets or class-level code that is evaluated against known examples. Benchmarks like Defects4J (Just et al., 2014) pose challenges for tools that require (correct) changes to a larger project, and are conceptually closer to experiences of developers. The recently proposed SWE-bench (Jimenez et al., 2024) offers a new angle on repository-scale changes by introducing natural language issues for program repair instead of a known test-suite. However, due to the natural language, some issues were too ambigious or un-solvable — spurring efforts of OpenAI to create a human-verified subset, called SWE-bench-verified (OpenAI, 2024a). Solving such natural language issues automatically brings the vision of a future AI software engineer closer to the capabilities of a human software engineer.
In this paper, we propose USEbench, a unified software engineering benchmark as a meta-benchmark which moves beyond project-level code editing tasks suggested by SWE-bench. Our goal is to unify existing individual benchmark sets and design a comprehensive unified dataset capturing multiple SE tasks. Our proposal combines a set of existing benchmark sets, namely SWE-bench-verified (OpenAI, 2024a), SWT (Mündler et al., 2024), REPOCOD (Liang et al., 2024) and REPOTEST. Out of these SWE-bench-verified captures tasks which involve code editing or program modifications to achieve bug fixing / feature addition. SWT bench presents validation tasks for testing real-world code fixes in software projects. REPOCOD is a recent benchmark which captures code generation tasks in a software project. REPOTEST is a not previously published derivative of REPOCOD: While REPOCOD encompasses repository level code-generation after removing a method body; REPOTEST removes tests for a specified method-body and evaluation is based on code-coverage.
The exact composition of USEbench from the constituent benchmarks shown in Table 1. Six different task types are covered, and for each task type several instances are included in USEbench. As a notable technical contribution, we provide a unified interface for agentic interaction with the unified benchmark. Previous benchmarks like SWE-bench provide a suitable interface for evaluation of code editing tasks i.e., after providing a .json file tasks are streamlined; still there is significant work involved in interacting with a software project, or extracting data from the constituent files of a software project. Thus, LLM agents proposed to work on SWE-bench tasks need to make these adjustments to work with SWE-bench, e.g. see (Orwall and contributors, 2025). We would want our unified benchmark USEbench to possess a greater degree of usability, so that LLM agents can be evaluated and compared in terms of their efficacy in conducting the software engineering tasks in the benchmark. To alleviate the burden on researchers, the constituent benchmarks are conveniently plugged behind a common interface revolving around docker images, providing options to read files and execute commands, regardless of the task. Due to the unified nature of our benchmark USEbench, we can also imagine unified software engineering agent (USEagent) working on the tasks in USEbench. Such a unified agent will be able to achieve complex software engineering tasks, which are solved via a sequence of the software engineering task types covered in the benchmark USEbench.
We now show two concrete scenarios of more complex software engineering tasks constructed from primitive task types in USEbench, namely incomplete fix and feature development, elaborated in Table 1. To concretely see how the different task types in USEbench are combined into compound tasks, let us consider the scenario of handling an incomplete code fix. In such a scenario - we (i) initially test the code, find failing tests, (ii) then generate a fix for handling the failed tests (iii) test the fixed code and find more failed tests (iv) finally generate a fixed code which passes all tests. These four steps could be four different runs of a unified software engineering agent USEagent, and could combine different task types from USEbench, namely code fixing and test generation. The unified API of tasks in USEbench allows us to combine and alter tasks to capture more complex software engineering challenges, like incomplete fixes.
Another scenario of a complex software engineering task, and focus of both industry and research (Wang et al., 2024b), is the complete addition of a feature to a code-base. This includes a problem description, and requires a functional method-body as well as accommodating tests. How to solve the task, as well as any form of decomposition, re-iteration or co-evolution is left up to the users of the benchmark. The final evaluation of this derived scenario will be against a hidden test-suite, as well as the generated tests.
The GitHub repository of USEbench will be made publicly available soon.
3. Background: Agents for Software Engineering
Agents typically utilize LLMs for decision making and content generation, along with autonomously invoked tools to interact with external entities. These tools allow the agent access to knowledge beyond the LLMs’ training data, and allow the agent to influence the external environment through the tool interface. Agents often employ reasoning frameworks to orchestrate the LLM and the tool usage. One of the popular reasoning frameworks used in agents is ReAct (Yao et al., 2023), in which the LLMs are instructed to reason about output-traces and invoke tools in an interleaved manner.
In software engineering, the target system that an agent interacts with is usually a software project. Another component of the external system is an execution environment, where software can be executed and tested. Existing software engineering agents (SE agents) proposed different ways to interact with the software project and the execution environment. AutoCodeRover (Zhang et al., 2024b) uses a set of project structure-aware tools (e.g. search_method_in_class to navigate the software codebase and gather necessary code context. It further performs test execution and spectrum-based localization (SBFL) (Wong et al., 2016) to pinpoint more relevant code locations. SWE-Agent (Yang et al., 2024) employs file-based tools to navigate files in the project, for example find_file, open, scroll_down, and together they form an agent-computer interface (Yang et al., 2024). OpenHands CodeActAgent (Wang et al., 2024b, a) provides a set of basic tools such as CmdRunAction to execute arbitrary bash commands inside a sandbox environment. A similar approach was followed by Google Jules agent (Google, 2024). Roughly summarized, the existing agents differ (intentionally) in how much structure and domain knowledge they introduce.
Another aspect of agent design is the decision framework. Existing SE agents have explored options for more controlled decision-making beyond the ReAct framework, in the context of software engineering. AutoCodeRover (Zhang et al., 2024b) and Agentless (Xia et al., 2024) targets software maintenance tasks, and they break the agent’s execution into phases like context retrieval/localization and repair. Another example is RepairAgent (Bouzenia et al., 2024), which restricts available tools through a finite-state machine. The different designs in existing SE agents raise an important question – how much autonomy should be given to the agent during decision making? On one hand, agents can follow a fine-tuned workflow for the targeted software engineering (SE) task (e.g. in AutoCodeRover, Agentless, RepairAgent). In this design, agents operate in fixed phases (e.g. localization is the first phase for program repair), which makes them more cost-effective (Zhang et al., 2024b; Xia et al., 2024). With the task broken down into phases, it also becomes easier to leverage program analysis techniques in each of the phases. However, agents with this design have less autonomy, and as a result, they could not be directly applied to another kind of SE task. Applying a program repair agent to test generation would require drastic changes to its workflow. On the other hand, agents can be given more autonomy by having access to a general set of tools and the freedom to decide when to use these tools (e.g. in SWE-Agent and OpenHands CodeActAgent). This general design makes the agents applicable to many kinds of software engineering tasks. However, this generality makes it harder to exploit domain knowledge and program analysis in the agents. Moreover, when the task becomes complex and the agent execution gets longer, it is challenging to interpret the agent trajectory consisting of a list of general tool usages. In comparison, agents operating in “phases” are more interpretable, since we can inspect the results of each phase (e.g. what is the identified faulty location at the end of a localization phase).
In this work, our goal is to have an agent applicable to multiple types of software engineering task, while still utilizing domain-specific optimizations and interpretability. We name this agent Unified Software Engineering Agent (USEagent).
4. Unified Software Engineering Agent
The first step towards a unified software engineering agent is breaking down the fixed workflow of existing agents into components (i.e., phases) and add orchestration to decide which component to invoke based on the task type and task state. To achieve this, we propose a Meta-Agent, a central LLM-reasoning agent that orchestrates various components, henceforth called actions. The Meta-Agent and the actions together form a unified software engineering agent (USEagent) that is capable of solving several types of software engineering tasks with a higher level of autonomy. We identify several challenges in designing such a unified agent to solve tasks in USEbench.
The first challenge is to adapt to tasks beyond a fixed workflow. Previous agents extract actions such as context retrieval or patch generation as clear “units of work” from developer workflows and organize them into state-machines or fixed workflows. This rigid organization of actions restricts their adaptability to other tasks. To achieve high agent autonomy we allow the actions to be freely composed and introduce a Meta-Agent that orchestrates the actions using a ReAct-style loop, forming a meta-layer over the actions, shown in Figure 1.
Figure 1. Concept: Meta-Agent abstracting over actions
Starting from natural-language descriptions, the Meta-Agent invokes actions, observes the changes made by the action, reasons about the output, and decides on the next action to take. This framework enables on-the-fly composition of workflows, adjusting to different tasks, errors and changing states. ReAct-style decision making has been widely used in previous agents to orchestrate tools that directly interact with the codebase; which we extend towards action orchestration. Subsequent to the autonomy is the question of termination: A USEagent must be able to terminate ad libitum when the task is considered completed. We suggest a termination action, that the Meta-Agent can choose to invoke. The overall trajectory and choices of the Meta-Agent will construct an action-graph. In our action-graph, any node can be the start or end-point - for example, depending on the task, an action of “executing tests” can be the starting point or the last assurance before termination. The agent execution can be seen as an implicit construction of this action-graph on the fly.
The next challenge is to provide the Meta-Agent with a set of reliable, yet flexible actions that are suitable for composition. Existing general agents such as OpenHands CodeActAgent (Wang et al., 2024b) and Google Jules (Google, 2024) provide the LLM with a console to execute any command. The action of executing a console command is extremely flexible, but hard to control. The agent’s execution consists of a sequence of console commands, which are difficult for human developers to interpret and raise (security) concerns. Thus, we propose to design our actions at a coarser granularity, where each action encapsulates a “unit of work” that developers carry out during the software engineering process. Each action receives instructions from the Meta-Agent on what outcomes it is supposed to achieve, rather than how it achieves them. The Meta-Agent and individual actions communicate through this intent-based interface. An example action is EditCode, where the Meta-Agent specifies what should be edited in the code, but not the exact details of making the edits. This design results in a set of modularized actions with clear responsibilities, that can be tested, extended and maintained individually. Moreover, when one existing action is improved, the improvements propagate across all involved task types. We present a sample list of actions for unified software engineering agents in Section 4.1.
The third challenge is knowledge management: For complex tasks, the information provided a priori is not sufficient, and a central effort in solving tasks is identifying and producing relevant information. Once produced, this context should be made available to the agentic system, i.e. Meta-Agent and other actions. Existing work has defined this problem as memory management of agents (Zhang et al., 2024a; Han et al., 2024), where memory refers to historical information relevant to the current task. In literature on LLM agents, memory is often categorized into short-term (e.g., information within an ongoing conversation), long-term (e.g., historical data from previous sessions), and consensus memory (e.g., shared knowledge between multiple agents) (Han et al., 2024). In this work, we explicitly identify what information to memorize when designing a unified software engineering agent, and classify them into the different types of memory. We categorize the results of actions as short-term memory, which is only presented to the Meta-Agent to decide the next action. Project-specific knowledge, such as developer-written documentation, is considered long-term memory. These documents are embedded and stored in a vector database accessible for Retrieval-Augmented Generation (RAG) by the agent. Finally, artifacts generated by actions that are useful in guiding other actions constitute consensus memory. Such artifacts include program locations or code edits that can be notable to other actions. We introduce a structured task state that represents the consensus memory, which can be read and modified by different actions.
4.1. Instantiating AutoCodeRover as USEagent
In this section, we highlight steps to build the unified SE agent USEagent that handles multiple types of SE tasks, from an existing SE agent that only focuses on one task type. We build our USEagent on top of the open-source agent AutoCodeRover (Zhang et al., 2024b; Ruan et al., 2024), since AutoCodeRover operates in different phases which are suitable as starting points for actions. We elaborate the changes to AutoCodeRover and flesh out our proposed solutions to the challenges in Section 4. Fine grained implementation details, prompts, etc. can be found in the supplementary material.
AutoCodeRover is an LLM agent that targets program maintenance tasks. Figure 2 shows the overall workflow of AutoCodeRover. Starting from a natural-language description of a software issue, AutoCodeRover first writes an executable self-contained test that reproduces the issue (Step ①). The stacktrace of the reproducer test, together with the task description, is sent to a context retrieval component to identify relevant code locations (Step ②). The relevant program locations are made available to the Edit Code component for writing a candidate patch (Step ④). With both LLM-generated patch and reproducer test, the Review Patch component then executes the tests on the patched program, decides whether the are correctly written, and iteratively improves them (Step ⑤). Finally, the AutoCodeRover workflow finishes when the reviewer approves a patch or the execution reaches a pre-defined round limit (Step ⑥). AutoCodeRover proved effective for resolving software issues such as bug fixing (Zhang et al., 2024b; Ruan et al., 2024). However, these fixed transitions among components make agents like AutoCodeRover not generalizable to task types that require a different workflow. For example, in a task like test-suite generation, steps such as Reproduction and Review Patch are irrelevant, requiring the design of a new workflow or even a new agent.
Figure 2. Overview of AutoCodeRover. AutoCodeRover composes a fixed workflow for program maintenance tasks.
Figure 3. Overview of the USEagent and workflow. The Meta Agent chooses available actions, provides the state and retrieves a altered state until termination is chosen.
Instead of designing a new agent for each task type, our goal is to build one general agent that handles multiple task types. To achieve this, we disassemble the workflow of AutoCodeRover and reassemble it into USEagent shown in Figure 3. We adapt components in AutoCodeRover such as Code Retrieval into actions that can be freely composed by a Meta-Agent for different task types. Given the task description that mentions the task type in natural language, the Meta-Agent uses ReAct-style reasoning to select next actions to execute. Short-term feedbacks such as action execution results are reflected directly to the Meta-Agent, while artifacts that should be preserved longer are stored into a task state representing the consensus memory among actions. We next discuss each component of USEagent in more details.
Action Description Action Signature Interaction with Task State
CodeRetrieval Collect program code relevant to the task description. Description $\rightarrow$ LocationsCode Write: LocationsCode
TestRetrieval Collect test code relevant to the task description. Description $\rightarrow$ LocationsTests Read: LocationsCode
Write: LocationsTest
Reproduction Generate a system-level standalone reproduction test. Description $\rightarrow$ DiffTest Write: TestReprod.
Write: Output(TestReprod.)
ExecuteTests Execute a subset of the program test suite List[TestFile] $\rightarrow$ Summary(Results) Read: LocationsTest
Write: CommandTest
EditCode Make modifications across the codebase, including test-code. $\Delta$Behavior $\times$ Location $\rightarrow$ Diff Read: LocationsCode
Read: LocationsTest
Write: Diff
ReviewPatch Review and iteratively improve a generated patch and generated test. TestReprod.$\times$×DiffCode $\rightarrow$ Bool Read: DiffCode
Read: TestReprod.
Write: DiffCode
Terminate Finish the agent execution and select a final solution. State $\rightarrow$ Diff -
Table 2. Details of actions used in USEagent. Diff denotes a Patch to the project, and subscript specifies modifiers.
E.g. DiffCode expresses a change to program code, instead of test code.
Actions.
Given the tasks of program repair, regression testing, code generation, and test generation in USEbench, we identify a priming set of actions for USEagent. The set of actions are shown in Table 2. Each action exposes its interface consisting of the description, input and output to the Meta-Agent, but does not expose execution details within the action. In addition to the input/output interaction with Meta-Agent, each action can also read/write to the task state.
Many of the actions in Table 2 come from existing agents like AutoCodeRover (Figure 2), and incorporating them into USEagent only involves specifying their input/output interface. On top of the existing agents, we designed a few new actions that are essential for the task types in USEbench. These new actions include TestRetrieval and ExecuteTests. The TestRetrieval action explores the codebase and retrieves testcases relevant to the current task. It works similarly compared to the existing CodeRetrieval, which employs a set of search tools such as search_func to search for relevant code units in the program. The ExecuteTests action executes parts of the project test suite, which is useful for validating code/test changes by execution. Current agents focusing on SWE-bench usually assume the commands to run existing unit tests are given as an input to the agent (Ruan et al., 2024; Xia et al., 2024); however, this assumption requires additional manual configuration when setting up the agents on new projects. Our ExecuteTests action does not assume the test commands to be given, and instead it queries the project documentation files to retrieve the project-specific commands through RAG. In addition, we added a special Finish action that signals the termination of the agent execution. The Finish action is invoked with an argument specifying the final result for the current task. Unlike other actions which encapsulates concrete workflows, the Finish action only terminates the agent and outputs its argument as the final result from the agent.
These actions can be invoked multiple times and combined in various ways to complete a given task. For example, for a code generation task, the Meta-Agent may first invoke CodeRetrieval to understand the surrounding code context, invoke EditCode to draft the implementation, invoke TestRetrieval to find out where existing tests for this code component resides, and then invoke EditCode to write new tests for the newly added code. It may then enter a refinement loop of improving the new code and tests using the ExecuteTests and EditCode actions. We note that this sequence of action invocation can happen without prior configuration, as we will show in section 6.3.
Task State.
Our next contribution towards a USEagent is the design of a task state. The task state represents the consensus memory among the actions - each action can write new artifacts generated by them into the task state so that the artifacts can be visible to other actions. Our task state captures essential artifacts involved in a software debugging process, including relevant locations, test execution information, and prior attempts of code modifications. Specifically, the task state $S$ is defined as: $S=(L_{c},L_{t},R_{exec},DS)$. Here, $L_{c}$ represents the relevant code locations (e.g., program methods and classes); $L_{t}$ represents the test locations (e.g., unit test methods). The locations are identified by the retrieval actions, and are utilized by downstream code editing actions. $R_{exec}$ contains the execution results of different tests against different candidate patches. These execution information provides guidance to other actions in improving partial solutions and selecting the final solution. The last part of the task state is a diff store ($DS$), which records all the generated diff contents from the EditCode, Reproduction and ReviewPatch actions. These diff contents include modifications to both the program code and the test code. The diff store is necessary to allow for versatile invocations of the ExecuteTests and EditCode actions. By selecting different subsets of diffs in $DS$ and passing them to ExecuteTests, tests can be executed on different versions of the project after applying the selected diffs. Similarly, EditCode can first apply an existing patch from $DS$, and subsequently introduce additional modifications, thereby enhancing partial patches.
MetaAgent.
The last important addition is the Meta-Agent which orchestrates the various actions. As shown in Figure 3, Meta-Agent takes in the task description and an initial empty task state (Step ①), and then iteratively selects an action from the action space to execute (Step ②). Each action represents an encapsulated workflow to complete a “unit of work”, and exposes its input/output interface to the Meta-Agent. At each step of action selection, we present the Meta-Agent with the current task state, the overall task description, as well as the output from the previously invoked action. The invoked action modifies the task state and retunrs it back to the Meta-Agent (Step ⑤). This interleaving of reasoning (i.e. select the next action based on current output and state) and action (i.e. execute the actual action to obtain new observations) forms a ReAct-style loop at the action level. When the Meta-Agent deems that one of the diffs in the diff store is a satisfactory solution to the given task, it invokes the Finish action to end the agent execution and output the final solution. In the case where the Meta-Agent could not decide on the final solution (i.e. could not invoke the Finish action) after a pre-defined limit for the ReAct-loop, we ends the execution and invoke the LLM to select one of the candidates in the diff store as the final solution.
4.2. Configuring OpenHands
OpenHands CodeActAgent (Wang et al., 2024b) is a general-purpose agent designed to solve general tasks [^1] Note: OpenHands is also developing CodeActSWEAgent, specialised on SWE-Bench. We purposefully chose the non-specialised implementation for USEbench. The entries on the SWE-Bench leaderboards are to our knowledge results from CodeActSWEAgent.. It utilizes a structured controller-agent-runtime architecture, where the AgentController serves as a supervisor that enforces operational constraints (conversation iterations, budget) and manages the agent’s lifecycle (start/stop/pause), while the CodeActAgent itself makes decisions, interpreting LLM responses and converting them into actions to interact with the runtime, which are isolated sandbox environment. Since CodeActAgent does not rely on a predefined task-specific workflow, it can be applied to a wide range of software engineering tasks, including those in USEbench, without requiring modifications to its architecture.
During execution, OpenHands CodeActAgent operates iteratively, generating actions, executing them, and processing observations before determining the next step. The actions include:
  • CmdRunAction: execute Linux bash commands.
  • IPythonRunCellAction: execute Python code in Jupyter or IPython environment.
  • FileEditAction: read, write, or edit files in the runtime.
  • AgentFinishAction: signal task completion or termination.
The execution of CodeActAgent terminates when the agent issues an AgentFinishAction, reaches resource limits (e.g., maximum iterations), or encounters an error. In practice, we design different prompt templates for each task type as input to OpenHands CodeActAgent, incorporating both the task type and task description. Compared to USEagent, the OpenHands CodeActAgent employs a similar reasoning framework, but the actions operate on a lower-level such as execution of a single bash command. It follows a single ReAct-loop centered around a more open set of available actions. This free-wheeling approach is an alternative to the more structured and pronounced options in Section 4.1.
5. Research Questions & Experiment Setup
Our first research question focuses on evaluating and comparing agentic systems for tasks beyond program repair.
RQ1: Efficacy of Agentic Systems How well do state of the art agentic systems work when facing varying software engineering tasks from USEbench?
We answer RQ1 by adapting AutoCodeRover and OpenHands before applying them to all data points outlined in Section 2 and report a PASS@1. In addition to the efficacy, we also report an investigation of the major obstacles for the individual benchmarks. To gather information on retrys and estimate the effect of randomness, we perform a PASS@5 on a significant subset of the datapoints.
A known issue and motivation for our next research question is overfitting in plausible patches. Such patches pass the evaluation-criteria by bypassing the harness with tricks rather than benign functionality. Sometimes, actual solutions are checks for edge-cases - these can only be distinguished from overfitting by consulting the ground truth (i.e., the gold patch). We aim to identify (and quantify) false-positives and other cases of overfitting solutions.
RQ2: Analysis of Plausible Patches What is the rate of false positives among the resolved patches?
We sample a significant subset222at 95% confidence with 10% error margin of solved datapoints and manually examine whether the plausible solution is overfitting. During this manual investigation, we also look for memorization in plausible patches, i.e. patches that perfectly mirror the ground truth developer patch or contain suspicious content.
As described in Section 4, one milestone is to make the agent adapt autonomously to different tasks. We study this adaptation by investigating the resulting sequence of actions in RQ3:
RQ3: Effectiveness of self-configuration Can USEagent self-configure to different task types?
We present the sequence of actions of the USEagent and accumulate their patterns per task. We showcase examples on how these top-level patterns translate into commands, queries and edits.
Lastly, we identify open challenges in the current unified software engineering agent. We formulate key difficulties and discuss potential future directions to address them.
RQ4: Open challenges for agentic systems What are the challenges faced by the current unified agent in tasks in USEagent? Can we identify actions to remediate them?
Experiment setup.
We evaluate agentic systems such as USEagent and OpenHands CodeActAgent on the full USEbench which consists of 1271 datapoints. In our experiments, all agentic systems use Anthropic Claude 3.5 Sonnet v2 (claude-3-5-sonnet-20241022) as the backend LLM. For USEagent, we set the maximum rounds of action invocation from the Meta-Agent to 20, and force the agent to end execution and select a candidate solution. We set the model temperature to zero in USEagent. To better understand the effect of randomness, we randomly sampled a statistically significant subset (295 instances) to report PASS@5 results from both agentic systems.
6. Results
6.1. RQ1: Efficacy of Agentic Systems
An overview of the efficacy of the agents on USEbench are shown in  Table 3. We investigate the results per system.
Table 3. Efficacy of USEagent and OpenHands CodeActAgent on USEbench.
USEagent OpenHands
Dataset Tasks PASS@1 PASS@5 PASS@1 PASS@5
SWE-Ver 500 228 (45.6%) 66.7% 192 (38.4%) 52.1%
SWT 298 120 (40.3%) 53.6% 85 (28.4%) 50.7%
REPOCOD 200 12 (6.0%) 15.2% 11 (5.5%) 6.5%
REPOTEST 173 55 (31.8%) 42.5% 45 (26.0%) 42.5%
SWETRY 100 8 (8.0%) 30.4% 7 (7.0%) 8.7%
Total 1271 423 (33.3%) 49.5% 340 (26.8%) 44.1%
USEagent.
[…]
While being effective on issue resolution tasks, USEagent can handle other tasks which specialized agents could not be applied to. For issue reproduction tasks in SWT, USEagent resolved 40.3% of the tasks. Here, “resolved” means USEagent generated test cases that cover all changed lines in the developer-written patch for the issue, without seeing the patch. Moreover, for a significant number of unresolved tasks, USEagent generated test cases that achieve a coverage above 90%. In these cases, the generated tests do not fully satisfy the requirement of the benchmark, but only miss coverage on individual statements or branches. On test generation tasks in REPOTEST, USEagent achieves efficacy of 31.8%, which is slightly lower than similar tasks in SWT. While the requirement in SWT is to cover a developer-written patch, the requirement in REPOTEST is to generate tests that cover an entire method, which can be more challenging. There are generated tests that cover most scenarios in a method, but do not cover all the lines. For example, one pattern in the unresolved tasks is that the agent provides only one positive test-input that executes the “normal” paths of a method, followed by tests that raise exceptions. In contrast, the ground-truth tests written by developers not only focus on exception paths, but extensively test the “normal” inputs. We argue for certain advantages in agent-generated tests over the developer-written tests; the agent often generates shorter tests where each test has individual assertions, while the developer-written tests tend to consist of long tests that combine multiple inputs and assertions. In a real test-suite, short individual tests can provide more informative error messages when part of the test-suite fail (Spadini et al., 2018; Van Deursen et al., 2001).
Tasks in REPOCOD proved to be the most challenging with only 6% resolution rate. REPOCOD tasks are generally challenging because the requirement is to generate a complete method implementation that can pass a large number of hidden test cases. Overall, for many unresolved instances, the generated method passes a high number of tests, yet does not account for a few edge-cases. The majority of resolved instances stem from Sympy, a library for symbolic mathematics. We expect a relation between the relative success in Sympy and the fact that mathematical reasoning is becoming a standard element of training data in foundational models.
On the compound benchmark SWETRY, USEagent has a resolve rate of 8%, which shows some promise to apply USEagent as a follow up to partial fixes. We need to stress that the datapoints in SWETRY are sourced from previous failed attempts of agentic systems, which means they are from the more challenging instances in SWE-verified. We observe two main failure modes on the SWETRY tasks. Firstly, since the partial patch is given to the agent as part of the task description, the agent often attempts to derive a solution on top of this partial patches by making modifications to it. However, the partial patch may be misplaced. In these cases, the agent still chose to iterate over this patch, and failed to discard the partial attempt and approach the task in a different way. Secondly, the Meta-Agent sometimes ends the agent execution prematurely. For example, when the ExecuteTests action still shows some relevant test errors, the Meta-Agent wrongly disregards errors as irrelevant to the task, ending the workflow. This can be addressed by configuring the prompt, at the cost of longer execution time and potential overfitting in other benchmarks.
OpenHands.
OpenHands CodeActAgent achieves a PASS@1 of 38.4% on SWE-verified, which is lower than the reported result from the SWE-Bench leaderboard (SWE, [n. d.]). We attribute this difference to the choice of agent - the reported result on the leaderboard is generated using a specialized CodeActSWEAgent, while we used the more general CodeActAgent in our evaluation.
We utilize the general CodeActAgent to solve the different types of tasks in USEbench. For SWT, CodeActAgent achieves a resolution rate of a 28.4%, slightly lagging behind USEagent. Similar to that of USEagent, CodeActAgent only achieves a low resolution rate on REPOCOD (5.5%). However, unlike USEagent, resolved instances are evenly distributed across sympy, astropy (a popular library for Astronomy and Astrophysics), and Plotly (a graphing library), each contributing 27% of the total resolved cases. On REPOTEST, OpenHands CodeActAgent resolves 26.0% of the total instances, which is similar but slightly lower compared to its performance on SWT. This trend is consistent with that of USEagent. On SWETRY, CodeActAgent only solves 7%, similar to that of USEagent, indicating that improving partial fixes is still a difficult task for general agents.
PASS@5
Moving to a PASS@5 increases the efficacy of USEagent to 49.5% (+48.6%) and OpenHands to 44.1% (+64.6%). This shows a similar increase across the two systems when employing global retries. Both systems largely benefit when facing edge-cases, e.g. to cover the last statement in testing or finding a correct reproducer.
Summary: RQ1 - Efficacy of Agentic Systems USEagent solves the tasks from USEbench at a 33.3% rate, compared to 26.8% of OpenHands. There is a consistent improvement over all benchmarks. USEagent’s efficacy of 45.6% for SWE-verified is close to the performance of fine-tuned systems.
6.2. RQ2: Dissection of Plausible Patches
For USEagent, we manually investigated a sample of 218 plausible solutions (i.e. passing the resolution criteria of each benchmark) to understand the degree of memorization, overfitting or other anomalies. We report a low number of 3 cases of memorization (1.3%), next to 23 cases of overfitting (10.5%). The percentage of overfitting solutions is lower than the previously reported overfitting rate of 31% on SWE-bench from AutoCodeRover (Ruan et al., 2024). We attribute the lower overfitting rate in part to our extensive ExecuteTests action in USEagent, which allows for more versatile test execution. As a result, we see an overfitting rate of 13.6% from USEagent on the SWE-verified dataset. Another reason for the lower overfitting rate is from USEbench: the SWT and REPOTEST datasets show low numbers of overfitting, which further reduces the overall rate. Overfitting patches usually bypass tests using if-else for early return values. Sometimes, this matches the intended handling of edge-cases, whether such patches are overfitting was determined by distance to ground truth. Furthermore, we identified a set of 33 anomalies that are plausible, and do not overfit, yet they are suspiciously different from the ground truth. One example anomaly includes the introduction of a new vector library for xarray (a vector library itself). Another example is an unorthodox class-inheritance check for scikit-learn - instead of utilizing pythons standard function isinstance, a predicate is applied to the objects obj.__classes__. These solutions are functionally correct but are unlikely to be acceptable by developers, so we have marked them as anomalies.
Summary: RQ2 - Analysis of Plausible Solutions We manually inspected 218 plausible solutions from USEagent and report 10.5% overfitting, 1.3% memorization. Overfitting was most common within SWE-verified, and least common in REPOTEST and SWT.
6.3. RQ3: Self-Configuration
(a) Distribution of SWE Steps - the most common path starts with reproduction followed by code retrieval, editing and review.
(b) Distribution of SWT Steps - most common trajectory begins with test-case retrieval, test-editing and execution.
Figure 4. Comparison of SWT and SWE Step Distributions for USEagent on a significant subset of USEbench. Each bar shows how often different actions are invoked at a particular Meta-Agent step.
Figure 4 shows the different choices of the Meta-Agent when facing different benchmarks (we present two for brevity). The histograms display the actions taken first, second, etc. including the termination action Terminate. We observe some clear patterns in the different benchmarks that match our intuition. For program repair tasks (i.e. in SWE-Verified), we see in Figure 4(a) that the first step prominently consists of generating a reproduction test, followed by retrieving relevant code context. Over the course of agent execution, the EditCode and ExecuteTests become more prominent, introducing and verifying code changes to the software. For regression testing (in SWT), Figure 4(b) shows that by-large the first step chosen is TestRetrieval, immediately followed by a (test-)code edit. The majority of USEagent trajectories on SWT tasks then consists of alternating test-changes and test-executions, as shown in the histograms. Although we observe a pattern in the sequence of action invocations, some invocations seem noisy. For example, Reproduction action is to generate a system-level reproducer test, but it is still invoked in some of the SWT trajectories (the requirement from SWT is to generate unit-level tests). We manually investigated some cases and found that the reproducers were later transformed into unit-level tests. Nevertheless, the sequence of action invocation follows a certain pattern for each task type as shown in Figure 4, showing that USEagent and self-configure its workflow based on the task type. Across all benchmarks, USEagent first gathers information (using the retrieval actions or reproducers), converging towards alternating EditCode and ExecuteTests until either budgets are exhausted or Terminate is chosen. This behavior showcases the capability of emulating concepts like Test-Time Scaling (Brown et al., 2024) and command-discovery (Bouzenia and Pradel, 2024) through the Meta-Agent.
Summary: RQ3 - Self Configuration USEagent is capable of choosing different action-patterns for different benchmarks: Program repair (SWE) starts with Reproduction, regression testing (SWT) with TestRetrieval. Initial steps generate information before the actions converge towards alternating EditCode and ExecuteTests until termination.
6.4. RQ4: Open Challenges in Agentic Systems
Based on our previous investigations, we identify several major challenges in the current unified software engineering agent. Firstly, when the task requires writing a large amount of code (e.g. feature development tasks), the agent often generates solutions that are “mostly correct”. This is reflected on our evaluation on the REPOCOD dataset, where the task is to implement a complex function based on natural language requirement, and the correctness is defined as passing a large number of hidden unit tests. We observe a number of solutions generated by the agent implement the required feature almost correctly, but fail a few hidden tests due to missing edge cases. The edge-case behaviors are sometimes not specified clearly in the natural-language task description. We attribute this challenge in part to the inherent ambiguity of natural language. A future agent may first resolve the ambiguity in natural language by transforming the task requirements into more “formal” artifacts such as specifications and test cases, and then generate solutions based on the formal specification. Another approach is to design human-agent interaction schemes to clarify ambiguity when needed.
Secondly, the current agent lacks the capability of “backtracking” when the agent execution does not yield meaningful results on a given execution path. For example, when there is a partial patch at a particular program location, the agent is prone to continuously make small edits to the partial patch and execute tests to verify them. However, the agent may not make good progress with this approach, since it may be impossible to craft a good solution at the location of the partial patch. This issue is amplified in the SWETRY dataset, in which the task description comes with a partial patch. To tackle this issue, a future agent should employ a backtracking mechanism to discard unpromising partial solutions, and start over at a previous step. It could be possible to employ the recent reasoning LLMs to examine the agent execution trace and decide whether to backtrack, since reasoning models have shown backtracking behaviors in their thinking process (Guo et al., 2025). Moreover, recent work (Antoniades et al., 2024) on employing search algorithm on agent execution trajectories can also help the agent to escape from unpromising paths.
Lastly, we still observe patch overfitting in tasks such as program repair and regression testing. Overfitting is a known problem for automated program repair (Goues et al., 2019). Overfitting happens due to the generated patches passing given tests, but missing actual requirements. Thus, to avoid overfitting, if the tests are too specific, we need to generalize them. Agentic systems show promise for test generation (also reflected in this work), and the framework introduced by USEagent allows for additional agents that cover test-amplification or mutation testing, adversarial reasoning or test carving. The produced artifacts will feed back into test-execution, code generation, review, etc. and reduce overfitting as well as increasing efficacy and trustworthiness.
Summary: RQ4 - Open Challenges Challenges consist of exploring edge-case handling and alternative solutions in feature development, backtracking and discarding poor partial results in repair, and repair-overfitting. We suggest search-approaches for the first, and agents for test-amplification for the latter.
7. Related Work
SE - Benchmarks
Recent Software Engineering Benchmarks fall into two broad categories: Isolated coding exercises and repository level tasks. Some coding exercises like CodeXGlue (Lu et al., 2021) come without an evaluation metric and form a training corpus, others like HumanEval (Chen et al., 2021a), MBPP (Austin et al., 2021), ClassEval (Du et al., 2023) or CoderEval (Yu et al., 2024a) provide a challenge accompanied by an evaluation harness (i.E. a test-suite). There are already ongoing efforts in providing a meta-benchmark for these tasks (Yu et al., 2024b)(Liu et al., 2024), which commonly target LLMs directly, less agentic systems. For USEBench, we select repository-level tasks to provide challenges of a uniform granularity. We decided against coding exercises as many models slowly saturate on them (Sallou et al., 2024).
Large Language Models and Agentic Systems
We consider a system agentic if the LLMs have a certain autonomy and are able to interact with the system through means of commands or other pre-defined tools. As such, Agentless (Xia et al., 2024) utilizes LLMs but is no agentic system due to its pre-defined workflow. Within agentic systems we see a divide into raw LLM-Agents that interface LLMs directly with a console and minimal tooling (e.g. ReAct (Yao et al., 2023), Gemini (Google, 2024) or SWE-Agent (Yang et al., 2024)) and approaches that introduce more domain knowledge as well as additional techniques. These additions range from specialised tools (like reproduction and fault localisation in AutoCodeRover (Zhang et al., 2024b)) to self-reflective capabilities (like the reviewer-agent in SpecRover (Ruan et al., 2024)). For this work, we have focused on AutoCodeRover (Zhang et al., 2024b) and OpenHands (Wang et al., 2024b), as these are open source projects with big resonance in the community. While the AIDER project (AIDER Team, 2024) has a well-suited scope (multiple SE tasks), its workflow relies on human prompting. We position AutoDev (Tufano et al., 2024) similar to AIDER, as the primary interface is a CLI for human interaction.
Agentic Systems
Passerine (Rondon et al., 2025) by Google proposes a program repair framework oriented at SWEagent (Yang et al., 2024) incorporating a similar structure to our meta-agent, providing a total of 5 pre-configured commands (excluding free commandline). Our efforts differ by scope as we target more than program repair, and employ a larger set of agents to multiple tasks. Similarly, Microsoft’s MASAI (Wadhwa et al., 2024), defines an action-space for their central agent similar to the meta-agent, limited to program repair. We differ from both works by providing more agents, structure the systems’ information and responses through a task-state, and apply the agentic system to more tasks than program repair. CodeR (Chen et al., 2024) has a meta-agent-component that pre-defines an execution plan, but does not adjust it dynamically.
8. Threats to validity
Construct Validity
Within USEbench, we approximate testing capabilities exploiting test-coverage. Test-coverage does not imply a semantic coverage of the code, and can be achieved by various hacks. In lack of a better oracle we have opted for test-coverage as it still captures many of the attributes we want to evaluate: The patches must be in a correct format, location and fit into the projects test-suite. Furthermore, upon inspection many of the REPOTEST datapoints contain branches that handle edge-case behavior with errors, for which a syntactic coverage equals semantic testing.
Internal Validity
Data Leakage It is possible that the LLMs have been exposed to Code subject to evaluation or even datapoints from the source-benchmarks(Sallou et al., 2024). This is particularly likely as many publications (Yang et al., 2024; Ruan et al., 2024; Xia et al., 2024) utilize the APIs by common providers that use user-data for training. Combating this is beyond our capabilities, but we aim to provide an overview of memorization by manually inspecting a sample of positives.
9. Discussion
Cost Analysis
While we consider that costs will play a diminishing role for agentic systems as foundation models and infrastructure advances, we still want summarize the accumulated costs over the experiments. USEagent costs were at an average of 3.65 USD, with notable differences between sub-benchmarks. Herein SWE scored lowest with 2.20 USD average, and REPOCOD highest at a mean of 7.66 USD. The high costs at REPOCOD originate from the iteration of patch-generation and test-execution. Within actions, EditCode was the dominant cost-driver, due to the (large) contexts provided through relevant code and tests. We calculate the Pearson correlation coefficient for the elapsed time and the induced costs, and find a strong correlation of $r=0.89$. This shows a strong correlation between computational (e.g. test execution) and model (e.g. reasoning) efforts.
Perspectives
We consider the approach of USEagent in combination with the meta-benchmark USEbench to be a promising avenue for any task that revolves around code as its primary artifact. Many tasks such as dependency management or project configuration can be incrementally tackled by extending USEbench and introducing more actions into USEagent. For USEagent to become truly an AI Software Engineer, it must address tasks such as requirements engineering, data visualization, deployment, code review or even a-b-testing. The efforts presented in this work are the second step in agentic development towards an AI software engineer, the first being individual agentic systems for single tasks (like AutoCodeRover). With the USEagent we pave the way for a unified approach to interact autonomously with code regardless of the task, allowing us to extend our scope beyond its current limits. Going beyond the USEagent will involve studying the AI-AI and AI-human collaboration. Thus we will need to study the cooperative intelligence resulting from multiple USEagents and multiple human developers. This might establish the dynamics of future development teams.